OpenStudy (anonymous):

the half life of radon is 3.80days what would be its decay constant

OpenStudy (anonymous):

Well, if λ is the decay constant and you have exponential decay - do you know how the quantity N(t) at a given time t is defined? Do you understand, why we have exponential decay in case of radon?

OpenStudy (shayaan_mustafa):

\[\LARGE \frac{1}{2}=e ^{kt}\]take ln on both sides\[\large ln(0.5)=kt\]\[\large ln(0.5)=k(3.80)\]\[\large k=\frac{\ln(0.5)}{3.80}\]\[\large k=\frac{-0.693}{3.80}\]\[\large k=-0.18/day\]

OpenStudy (shayaan_mustafa):

Where k is decay constant and t is half life.

OpenStudy (anonymous):

Despite the fact, that this is wrong (the decay constant is obviously positive, when there is decay - hence the name).. Why do you just post the solution, @Shayaan_Mustafa ?? Yeah you want some super-cool medals, I get it.. but actually helping them get there is of much more value. This way, the questioner is just going to copy and not think about it..

OpenStudy (anonymous):

Half Life = \[T _{1/2}\] = \[\ln 2/k\] where k is the decay constant So k = \[\ln 2/3.8\] per day